OpenStudy (anonymous):

Two competing engineers have offered their solutions for the function. One has suggested g(x) = x3 – 2x2 + 4x – 8, and the other has suggested h(x) = x3 + 5x2 – 8x – 48. Explain to the C.E.O. of Oaville Lake Amusement Park why these other suggestions are invalid. Use complete sentences.

OpenStudy (anonymous):

please help!! I am very confused.

OpenStudy (anonymous):

whats the function, is this the full question

OpenStudy (anonymous):

this is the entire question. @dav21

OpenStudy (anonymous):

@TheRealMeeeee can you help me????

OpenStudy (therealmeeeee):

x has to be -12 or -13 so you don't go through the building if you pick x=-13, that looks safe so let's use x= 5, -4 and -13 a third-degree polynomial crosses the x-axis 3 times you can say y = (x-5)(x - -4)(x- -13) or y= (x-5)(x+4)(x+13) that will go through your points. It's y intercept is when x=0 or -5*14*13 = -910 that is very big. so let's multiply by 1/(14*13) = 1/182 that will make the y -intercept = -5 there ya go

OpenStudy (therealmeeeee):

there ya go goodnight